# Fluctuations of local plastic strain in granular media

### v1

@@ III.3 Quadratic terms from Bagi cells and DAWS
The quadratic strain terms obtained from the two methods are plotted in Figure 6 a) for the interval $[a_{k_{r}},a_{k}]$ and $\theta_{m}=16.8^{\circ}$ at position $P1$. The quantities $\langle\overline{\epsilon_{xx}^{2}}\rangle$, $\langle\overline{\epsilon_{yy}^{2}}\rangle$, $\langle\overline{\epsilon_{xy}^{2}}\rangle$, and $\langle\overline{\epsilon_{v}^{2}}\rangle$ in Bagi cells similarly increase with the number of oscillations and are all of the same order of magnitude. Their sum, $\langle\overline{\epsilon^{2}}\rangle_{o}=\left(\langle\epsilon_{v}^{2}\rangle+2\times\sum_{i,j}\langle\epsilon_{i,j}^{2}\rangle\right)$ ($i,j\neq z$), can be compared to $\langle\overline{\epsilon^{2}}\rangle_{u}$, which is derived from the inversion of Eq. 4 and includes all squared components of the strain tensor. Figure 6 a) shows that $\langle\overline{\epsilon^{2}}\rangle_{o}$ and $\langle\overline{\epsilon^{2}}\rangle_{u}$ are in good agreement. Therefore, comparing our experimental results with the theoretical predictions from [25] (Eq. 4), suggests that the quadratic terms $\langle\overline{\epsilon_{xz}^{2}}\rangle$, $\langle\overline{\epsilon_{yz}^{2}}\rangle$, and $\langle\overline{\epsilon_{zz}^{2}}\rangle$ can be neglected.
A comparison between both methods is also possible via the estimate of the root-mean-square relative displacements between grains. From the imaging method, computing the relative displacement $\Delta r$ between nearest neighbors yields the distributions $\Delta r_{x}$ and $\Delta r_{y}$, as shown in Figure 6b). We find $\langle\overline{\Delta r_{x}^{2}}\rangle_{o}^{1/2}\approx 34~{}\mu$m and $\langle\overline{\Delta r_{y}^{2}}\rangle_{o}^{1/2}\approx 31~{}\mu$m. Turning to DAWS, we calculate from Eq. 3 and 4 that in all cases $V(\Delta\Phi)\gg\langle\Delta\Phi\rangle^{2}$. Hence, $V(\Delta\Phi)\approx\langle\Delta\Phi^{2}\rangle\sim q^{2}\langle\Delta r^{2}\rangle_{u}/3$ [15]. From Figure 6a), we find that $\langle\overline{\epsilon^{2}}\rangle_{u}\sim 10^{-3}$, which leads to $\langle\overline{\Delta r^{2}}\rangle_{u}^{1/2}\sim 47~{}\mu$m, similar to the value obtained from the optical method. It is important to note that the relative mean squared displacement $\langle\Delta r^{2}\rangle_{u}$ measured by multiply scattered waves does not depend on the symmetric strain tensor and is therefore distinct from the more commonly used mean square displacement derived from the displacement gradient tensor.
@@
III.4 Strain fluctuations : Intracycle diffusion and plasticity.
Since all measured quadratic strain terms are equivalent, and for the sake of conciseness, the remainder of the paper focuses on describing one specific strain component: the shear component (optical method). In particular, we emphasize how the shear strain $\epsilon_{xy,p}^{2}$ between $[a_{k},a_{k+1}]$ is derived from the consecutive forward (maximum) strain $\epsilon_{xy,m}$ over the interval $[c_{k},a_{k+1}]$ and the backward (return) strain $\epsilon_{xy,r}$ over $[a_{k},c_{k}]$, as illustrated in Fig. 4b). Data from the optical method are presented in the main text, while complementary data from DAWS, leading to the same conclusions and suggesting that the observed effects are not confined to the possible peculiarities of grains at the solid walls, are reported in Appendix C.
The variation of the quadratic strain as a function of both $\langle\overline{\epsilon_{xy,m}}\rangle$ and $\langle\overline{\epsilon_{xy,r}}\rangle$ from Bagi cells is plotted in Figure 7a) for all three zones of interest and the four oscillation angles. All data collapse onto a single curve, which can be approximated linearly by $\langle\overline{\epsilon_{xy,r,m}^{2}}\rangle=D_{o}\times|\langle\overline{\epsilon_{xy,r,m}}\rangle|$, with $D_{o}\approx 0.025$. The data exhibit consistent ”diffusive” behavior, driven by mechanical energy [28, 29, 30]. It is important to note that the term ”diffusion” is used here to describe the proportionality between strain fluctuations and mean strain, as seen in athermal amorphous systems [30]. The plasticity observed at the end of an oscillation arises from two consecutive forward and backward diffusive branches.
@@
Figure 7: a) Mean square strain from Bagi cells as a function of the average shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$ and the average backward shear strain $\langle\overline{\epsilon_{xy,r}}\rangle$. b) Plastic strain fluctuations from Bagi cells as a function of the maximum strain $|\langle\overline{\epsilon_{xy,m}}\rangle|$. In panels a) and b), circles correspond to position $P1$, and diamonds to position $P3$. Position $P2$, represented by stars, is plotted only in panel a). The colors red, blue, green, and violet correspond to measurements at $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$, respectively. Inset of b): Circles represent $A_{o}(\theta_{m})$ from Eq. 5 (Bagi cells), normalized by $A_{o}(10.8^{\circ})=3.3$. Triangles represent $A_{u}(\theta_{m})/A_{u}(10.8^{\circ})$ from DAWS (see Figure 11b in Appendix A), with $A_{u}(10.8^{\circ})=8.3$. The line is a linear fit to the data: $A_{o,u}(\theta_{m})/A_{o,u}(10.8^{\circ})=1.6\times\left(1-\frac{\theta_{m}}{\theta_{r}}\right)$.
The fluctuations of the plastic strain $\langle\overline{\epsilon_{xy,p}^{2}}\rangle^{1/2}$ from Bagi cells are plotted in Figure 7 b) as a function of $|\langle\overline{\epsilon_{xy,m}}\rangle|$ for zones of interest P1111 and P3333 and all the oscillations angles. Plastic quadratic terms from P2222 are not discussed because the values are too close to the noise level $\sim 10^{-4}$. A first observation is that $\langle\overline{\epsilon_{xy,p}^{2}}\rangle^{1/2}$ has a linear dependence on the mean strain, as opposed to the diffusive forward and backward strains above. Second, at fixed strain, higher oscillation angles give rise to lower fluctuations. This can be described by
$\delta\epsilon_{xy,p}\approx\langle\overline{\epsilon_{xy,p}^{2}}\rangle^{1/2}=A_{o}|\langle\overline{\epsilon_{xy,m}}\rangle|$ (5)
$A_{o}$ is a function of $\theta_{m}$ is plotted in Figure 7 b) (inset) showing a linear decrease. As shown in the Figure 11 of Appendix A we find similar behavior with DAWS. At this stage, our experimental data from two different methods show that (i) the strain is diffusive (ii) the fluctuations of the plastic strain vary linearly with the maximum shear strain and (iii) at fixed strain, fluctuations decrease with the oscillation angle.
@@ IV Discussion
To estimate the packing fraction $\phi$ of the material located in the zones $P1$ and $P3$, we use the relation $\phi\approx\phi_{0}\left(1+\langle\overline{s_{y}}\rangle/h\right)$, where $\langle\overline{s_{y}}\rangle$ is the position of the free surface, as shown in Figure 3 a). The difference in packing fraction between two systems sheared at different $\theta_{m}$ values is given by $\Delta\phi/\phi_{0}\approx\langle\overline{s_{y}}\rangle/h\ll\Delta\theta_{m}/\theta_{m}$. Moreover, at a fixed $\theta_{m}$, the relative variation of packing fraction during compaction is small, on the order of $\sim 10^{-2}$. Therefore, $\theta_{m}$ determines the constant maximum stress applied at each cycle, independent of $\phi$. However, the repeated oscillations at $\theta_{m}$ do modify the packing fraction by playing the role of mechanical taps, as discussed earlier.
We assume, as suggested by the shear cycles shown in Figure 4 a), a constitutive relation between $\epsilon_{xy}$, $\theta$, and $\phi$, where $\epsilon_{xy}(\theta,\phi)$ is an increasing function of $\theta_{m}$ and a decreasing function of $\phi$. Thus, Eq. 5 becomes $\delta\epsilon_{xy,p}=A_{o}(\theta_{m})\times|\langle\overline{\epsilon_{xy,m}}\rangle|(\theta_{m},\phi)$. In combination with the second relation, $A_{o}(\theta_{m})\propto\left(1-\frac{\theta_{m}}{\theta_{r}}\right)$ from Figure 7 b) (inset), we can describe the data.
To do so, the fluctuations of the plastic strain are re-plotted in Figure 8 b) as a function of $\phi$, estimated using the expression above (the following analysis is not sensitive to small variations of $\phi_{0}$ and $h$, nor to the particular method used to calculate the packing fraction.) At fixed $\theta_{m}$, the fluctuations of plastic strain grow with $\epsilon$ and correspond to a decrease in packing fraction. At fixed $\epsilon$ (iso-strain dashed lines in Figure 8 b)), the plastic strain fluctuations increase with decreasing $\theta_{m}$ via $A_{o}$, as observed in Figure 7 b). This in turn forces $\phi$ to decrease. Hence, we observe larger plastic fluctuations in looser systems both at fixed strain and at fixed stress. The case at fixed $\phi$ is more complex. Increasing $\theta_{m}$ increases the strain but decreases $A_{o}$. Comparison of iso-$\phi$ data at $\phi\sim 0.65$ in Figure 8 b) for $\theta_{m}=19.8^{\circ}$ and $21.8^{\circ}$ suggests that fluctuations increase with higher stress, and the stress-strain constitutive relation dominates over the moderate effect of $\theta_{m}$ on $A_{o}$.
Although the causality between the high fluctuations and the low packing fraction is not established, our results suggest that the free volume [31] could play a role in the emergence of plastic fluctuations. Following [31], an estimate of the average free volume per particle is $v_{f}=V_{g}\frac{\phi_{c}-\phi}{\phi_{c}^{2}}$ (where $V_{g}$ is the volume of a grain and $\phi_{c}=0.74$ is the close packing fraction), and assuming that $v_{f}$ is homogeneously distributed around the particles, one can estimate a characteristic ”free length”, $l_{f}$, available for the particle to rearrange. For the conditions of Figure 6 b) (with $\phi\approx 0.64$), we obtain $l_{f}\approx 85$ $\mu$m, to be compared with the root-mean-square relative particle displacements $\langle\Delta r^{2}\rangle_{o}^{1/2}\approx 30$ $\mu$m. The correspondence between the two orders of magnitude suggests that higher free volume provides more room for exploration of the topological landscape, leading to prominent chaotic irreversible relative displacements between the grains and higher fluctuations.
Lastly, we explain how the experimental relationship in Eq. 5 arises from the cumulative effect of two consecutive ”diffusive” branches. To illustrate this, we propose a straightforward description of a strain cycle, where the total plastic strain at the end of the cycle is the sum of the forward strain, $\langle\overline{\epsilon_{xy,m}}\rangle$, and the backward strain, $\langle\overline{\epsilon_{xy,r}}\rangle$ (see Figure 4 b)). As shown earlier, the mean square of the strain exhibits diffusive behavior, described by $\langle\overline{\epsilon_{xy,m,r}^{2}}\rangle=D_{o}|\langle\overline{\epsilon_{xy,m,r}}\rangle|$ where $D_{o}$ is the diffusion coefficient. Treating the strain in the Bagi cell as a random variable, the total strain can be written as $\epsilon_{p}=\epsilon_{m}+\epsilon_{r}$. The variance of the total strain is then given by $V(\epsilon_{p})=V(\epsilon_{m})+V(\epsilon_{r})+2\text{cov}(\epsilon_{m},\epsilon_{r})$, where $V$ represents the variance and cov is the covariance. We further assume that the forward and backward mean strains are related by a linear relationship $\langle\overline{\epsilon_{xy,r}}\rangle=-b\langle\overline{\epsilon_{xy,m}}\rangle$ where $b=b(\theta_{m})$ and $0\leq b\leq 1$. The parameter $b\to 1$ (or $b\to 0$) corresponds to a more elastic (or plastic) behavior. Consequently, the mean square of the plastic strain can be approximated as $\langle\overline{\epsilon_{xy,p}^{2}}\rangle\approx\alpha|\langle\overline{\epsilon_{xy,m}}\rangle|$ where $\alpha=D_{o}(1+b+2b^{1/2}r_{mr})$ and $r_{mr}$ is the correlation coefficient, defined as the covariance divided by the product of the standard deviations of the forward and backward strains, $\epsilon_{xy,m}$ and $\epsilon_{xy,r}$.
@@
The parameters $b$ and $r_{mr}$ can be determined experimentally. Fitting the sum of the variances of forward and backward strains as a function of $\langle\overline{\epsilon_{xy,m}}\rangle$ yields the parameter $b\approx 0.5,0.6,0.7$ for $\theta_{m}=16.8^{\circ},19.8^{\circ}~{}\text{and}~{}21.8^{\circ}$, respectively (not shown). Moreover, $r_{mr}$, which is plotted in Figure 8 a), is negative (indicating anticorrelation between maximum and backward shear) and varies linearly with the strain $r_{mr}=c|\langle\overline{\epsilon_{xy,m}}\rangle|-1$. The slope $c=c(\theta_{m})$, given in Figure 8 a), is a decreasing function. Arranging the various terms above gives $\alpha=D_{o}\left(A_{1}+A_{2}\times|\langle\overline{\epsilon_{xy,m}}\rangle|\right)$ with $A_{1}=(1+b-2b^{1/2})$ and $A_{2}=2b^{1/2}c$. $\alpha$ rationalizes our experimental findings in describing how diffusion, reversibility, and strain correlation couple to define plastic strain fluctuations.
Figure 8: a) Correlation between $\epsilon_{xy,m}$ and $\epsilon_{xy,r}$ strains in a shear cycle, with the slope $c$ indicated next to the linear fit of the data for each angle. Red, blue, green, and pink lines and points correspond to $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$ and $10.8^{\circ}$, respectively. b) Fluctuations of plastic strain as a function of the density of the packing estimated from the compaction near the free surface (see text). The crosses are the experimental points smoothed with a moving average of 4 points. Dashed lines represent iso-strain points, from bottom to top $\langle\overline{\epsilon_{xy,m}}\rangle=2.5\times 10^{-3},5.2\times 10^{-3},6.9\times 10^{-3}$. Plain lines are guides to the eye. The color code is the same as in a).
@@
With the expression for $\alpha$ above, fluctuations vanish as expected for $\epsilon\rightarrow 0$. However, for larger strains ($>10^{-3}$), the $A_{2}$ term dominates, and $\delta\epsilon_{xy,p}\sim\left(D_{o}A_{2}\right)^{1/2}|\langle\overline{\epsilon_{xy,m}}\rangle|$, which allows us to identify in Eq. 5 that $A_{o}=b^{1/4}\left(2cD_{o}\right)^{1/2}$. In our experiments, $D_{o}$ is constant, and $b^{1/4}$ varies slowly, so $A_{o}$ is primarily controlled by the parameter $c$. As a consistency check, $A_{o}$ is plotted as a function of $c^{1/2}$ in Figure 12 in the Appendix D. The linear fit $A_{o}=0.17\times c^{1/2}$ is in good agreement with the independent estimate $b^{1/4}\left(2D_{o}\right)^{1/2}\approx 0.2$. This identifies the different physical mechanisms that govern the empirical parameter $A_{o}$. Diffusion, due to frictional and hard-sphere interactions, enhances plastic strain fluctuations. However, these fluctuations are moderated by the high local strain anticorrelation in dense packings, which we tentatively attribute to more reversible trajectories, associated to the memory effect.
@@
To sum up, we find that the fluctuations of plastic strain are higher in the loosest packings, where the independence of forward and backward strains in Bagi cells is enhanced. Plasticity, viewed through the lens of local fluctuations, shares some common features but also exhibits differences with other divided systems such as emulsions and colloidal glasses. As in emulsions [10], fluctuations increase with decreasing packing fraction at fixed strain. However, in colloidal glasses [11], low packing fraction leads to fewer plastic events, as the lubricated interactions between grains allow for reversibility. Our study suggests distinct signatures of emerging plasticity in emulsions, colloidal glasses, and granular materials, shedding light on the effects of particle interactions.

### v2

@@ III.3 Quadratic terms from Bagi cells and DAWS
The cumulative plastic quadratic strain terms obtained from the two methods are plotted in Figure 6 a) for the interval $a_{k_{r}},a_{k}]$ and $\theta_{m}=16.8^{\circ}$ at position $P1$. The quantities $\langle\overline{\epsilon_{xx,p}^{2}}\rangle$, $\langle\overline{\epsilon_{yy,p}^{2}}\rangle$, $\langle\overline{\epsilon_{xy,p}^{2}}\rangle$, and $\langle\overline{\epsilon_{v,p}^{2}}\rangle$ in Bagi cells similarly increase with the number of oscillations and are all of the same order of magnitude. Their sum, $\langle\overline{\epsilon^{2}}\rangle_{o}=\left(\langle\epsilon_{v}^{2}\rangle+2\times\sum_{i,j}\langle\epsilon_{i,j}^{2}\rangle\right)$ ($i,j\neq z$), can be compared to $\langle\overline{\epsilon^{2}}\rangle_{u}$, which is derived from the inversion of Eq. 4 and includes all squared components of the strain tensor. Figure 6 a) shows that $\langle\overline{\epsilon^{2}}\rangle_{o}$ and $\langle\overline{\epsilon^{2}}\rangle_{u}$ are in good agreement. Therefore, comparing our experimental results with the theoretical predictions from [25] (Eq. 4), suggests that the quadratic terms $\langle\overline{\epsilon_{xz}^{2}}\rangle$, $\langle\overline{\epsilon_{yz}^{2}}\rangle$, and $\langle\overline{\epsilon_{zz}^{2}}\rangle$ can be neglected.
[…]
@@
III.4 Strain fluctuations : Intracycle variability and plasticity.
@@
Figure 7: a) Mean square strain from Bagi cells as a function of the average shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$ and the average backward shear strain $\langle\overline{\epsilon_{xy,r}}\rangle$. b) Plastic strain fluctuations from Bagi cells as a function of the maximum strain $|\langle\overline{\epsilon_{xy,m}}\rangle|$. Inset of b): Circles represent $A_{o}(\theta_{m})$ from Eq. 5 (Bagi cells), normalized by $A_{o}(10.8^{\circ})=3.3$. Triangles represent $A_{u}(\theta_{m})/A_{u}(10.8^{\circ})$ from DAWS (see Figure 11b in Appendix A), with $A_{u}(10.8^{\circ})=8.3$. The line is a linear fit to the data: $A_{o,u}(\theta_{m})/A_{o,u}(10.8^{\circ})=1.6\times\left(1-\frac{\theta_{m}}{\theta_{r}}\right)$. c) Mean strain for the backward branch of the shear cycles $\langle\epsilon_{xy,r}\rangle$ as a function of the forward branch $\langle\epsilon_{xy,m}\rangle$. d) Fluctuations of plastic strain as a function of the mean strain. Circles correspond to position $P1$, and diamonds to position $P3$. Position $P2$, represented by stars, is plotted only in panel a). The colors red, blue, green, and violet correspond to measurements at $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$, respectively.
Since all measured quadratic strain terms are equivalent, and for the sake of conciseness, the remainder of the paper focuses on describing one specific strain component: the shear component (optical method). In particular, we will show how the intracycle plastic shear strain fluctuations $\delta\epsilon_{xy,p}$ between $a_{k},a_{k+1}]$ emerge from the consecutive forward (maximum) strain $\epsilon_{xy,m}$ over the interval $a_{k},c_{k}]$ and the backward (return) strain $\epsilon_{xy,r}$ over $c_{k},a_{k+1}]$ (see sketch) in Fig. 4b). Data from the optical method are presented in the main text, while complementary data from DAWS, leading to the same conclusions and suggesting that the observed effects are not confined to the possible peculiarities of grains at the solid walls, are reported in Appendix C.
The variation of the quadratic strain as a function of both $\langle\overline{\epsilon_{xy,m}}\rangle$ and $\langle\overline{\epsilon_{xy,r}}\rangle$ from Bagi cells is plotted in Figure 7a) for all three zones of interest and the four oscillation angles. Although each mean strain corresponds to different states of compaction and applied stress, all data collapse onto a single curve, approximated by $\langle\overline{\epsilon_{xy,r,m}^{2}}\rangle=D_{o}\times|\langle\overline{\epsilon_{xy,r,m}}\rangle|^{\gamma}$, with $D_{o}\approx 0.025$ an effective diffusion coefficient and $\gamma\simeq 1$, consistent with random motion driven by shear [28, 29, 30]. We note, however, that data spanning several orders of magnitude in strain amplitude would be necessary to determine a more reliable estimate of the $\gamma$ exponent. Importantly, the term ”diffusion” is used here to describe the proportionality between strain fluctuations and mean strain.
Following the completion of each forward–backward shear cycle, the fluctuations of the intracycle plastic strain can be monitored across successive cycles, as shown in Fig. 7b, as a function of the maximum shear strain. Data from position P22 are not discussed because the values are too close to the noise level $\sim 10^{-4}$. First, for a fixed oscillation angle $\theta_{m}$, the fluctuations grow linearly with the mean maximum strain. Second, at a fixed maximum strain, larger oscillation angles result in smaller fluctuations. This can be described by
$\delta\epsilon_{xy,p}\approx\langle\overline{\epsilon_{xy,p}^{2}}\rangle^{1/2}\sim A_{o}|\langle\overline{\epsilon_{xy,m}}\rangle|$ (5)
$A_{o}$ is a function of $\theta_{m}$ and is plotted in Figure 7 b) (inset), corresponding to a linear decrease $A_{o}(\theta_{m})\propto\left(1-\frac{\theta_{m}}{\theta_{r}}\right)$. As shown in Figure 11 in Appendix C, we observe similar trends with DAWS. The parameter $A_{u}$, which is the DAWS equivalent of $A_{o}$, exhibits the same behavior as $A_{o}$ (Figure 7 b), inset).
With the constitutive relation $\epsilon_{xy}(\theta,\phi)$ (increasing function of $\theta_{m}$ and a decreasing function of $\phi$), Eq. 5 reads $\delta\epsilon_{xy,p}\sim A_{o}(\theta_{m})\times|\langle\overline{\epsilon_{xy,m}}\rangle|(\theta_{m},\phi(\theta_{m},k))$. It is therefore useful to clarify the combined effects of stress, strain and density in the emergence of plastic strain fluctuations. At fixed $\theta_{m}$, the packing fraction increases with $k$ so that $\epsilon_{xy}(\theta_{m},\phi(\theta_{m},k))$ decreases. From Eq. 5 fluctuations of plastic strain decrease as a simple effect since no plasticity is expected at vanishing forward shear amplitude. In order to examine the iso-density and iso-strain cases, $\delta\epsilon_{xy,p}$ is re-plotted in Figure 8 a) as a function of $\phi$, estimated using the expression in III.2.1 (the following analysis is not sensitive to small variations of $\phi_{0}$ and $h$, nor to the particular method used to calculate the packing fraction). Comparison of iso-$\phi$ data at $\phi\sim 0.65$ in Figure 8 a) for $\theta_{m}=19.8^{\circ}$ and $21.8^{\circ}$ suggests that fluctuations increase with higher stress, and the stress-strain constitutive relation dominates over the moderate effect of $\theta_{m}$ on $A_{o}$. Finally, at fixed $\epsilon_{xy}(\theta,\phi)$ (iso-strain dashed lines in Figure 8 a)), the plastic strain fluctuations increase with decreasing $\theta_{m}$ via $A_{o}$, as could be observed in Figure 7 b). This in turn forces $\phi$ to decrease. Thus, the higher iso-strain plastic fluctuations are associated with both lower packing fraction and lower stress level $\theta_{m}$.
Figure 7 c) shows that $\theta_{m}$ is key. The forward and backward mean strains are related by a linear relationship independent of the packing fraction $\langle\overline{\epsilon_{xy,r}}\rangle=-b\langle\overline{\epsilon_{xy,m}}\rangle$ where $b=b(\theta_{m})$ only, and $0\leq b\leq 1$. Parameter $b\to 1$ (or $b\to 0$) corresponds to a more elastic (or plastic) behavior. As determined in Figure 7 c), $b\approx 0.28,0.59,0.79$ and $0.87$ for $\theta_{m}=10.8^{\circ},16.8^{\circ},19.8^{\circ}~\text{and}~21.8^{\circ}$, respectively, showing that larger oscillation angles increase the reversibility of the mean strain. This behavior leads to Figure 7 d) where the variation of $\delta\epsilon_{xy,p}$ on its mean is established by rescaling the abscissa with $\langle\overline{\epsilon_{xy,p}}\rangle=(1-b)\langle\overline{\epsilon_{xy,m}}\rangle$. Remarkably, all data collapse onto a single master curve, revealing a unique dependence of the plastic fluctuations on the mean such that $\delta\epsilon_{xy,p}\sim C_{o}\left|\langle\overline{\epsilon_{xy,p}}\rangle\right|$, where $C_{o}\approx 5$ (see also Figure 11c) for the equivalent DAWS measurements).
Figure 8: a) Fluctuations of plastic strain as a function of the density of the packing estimated from the compaction near the free surface (see text). The crosses are the experimental points smoothed with a moving average of 4 points. Dashed lines represent iso-strain points, from bottom to top $\langle\overline{\epsilon_{xy,m}}\rangle=2.5\times 10^{-3},5.2\times 10^{-3},6.9\times 10^{-3}$. Plain lines are guides to the eye. b) Correlation between $\epsilon_{xy,m}$ and $\epsilon_{xy,r}$ strains in a shear cycle, with the slope $c$ indicated next to the linear fit of the data for each angle. Red, blue, green, and pink lines and points correspond to $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$ and $10.8^{\circ}$, respectively.
@@ IV Discussion
We now establish the link between the results presented in Figure 7. The plastic strain is $\epsilon_{p}=\epsilon_{m}+\epsilon_{r}$ so that treating the strain in the Bagi cell as a random variable constrains the variance to satisfy $V(\epsilon_{p})=V(\epsilon_{m})+V(\epsilon_{r})+2\text{cov}(\epsilon_{m},\epsilon_{r})$, where $V$ represents the variance and cov is the covariance. Along with $\langle\overline{\epsilon_{xy,m,r}^{2}}\rangle\sim D_{o}|\langle\overline{\epsilon_{xy,m,r}}\rangle|$, this provides $\langle\overline{\epsilon_{xy,p}^{2}}\rangle\approx\alpha|\langle\overline{\epsilon_{xy,m}}\rangle|$ where $\alpha=D_{o}(1+b+2b^{1/2}r_{mr})$ and $r_{mr}$ is the correlation coefficient, defined as the covariance divided by the product of the standard deviations of the forward and backward strains, $\epsilon_{xy,m}$ and $\epsilon_{xy,r}$. $r_{mr}$ can also be determined experimentally as in Figure 8 b). It is negative (indicating anticorrelation between maximum and backward shear) and varies linearly with the strain $r_{mr}=c|\langle\overline{\epsilon_{xy,m}}\rangle|-1$. The slope $c=c(\theta_{m})$, annotated in Figure 8 a), is a decreasing function. Arranging the various terms above gives $\alpha=D_{o}\left(1+b-2b^{1/2}+2b^{1/2}c\times|\langle\overline{\epsilon_{xy,m}}\rangle|\right)$.
@@
With this expression for $\alpha$, fluctuations vanish as expected for $\epsilon\rightarrow 0$. However, for larger strains ($10^{-3}$), $\delta\epsilon_{xy,p}\sim b^{1/4}\left(2cD_{o}\right)^{1/2}|\langle\overline{\epsilon_{xy,m}}\rangle|$, which directly yields the expression for $A_{o}=b^{1/4}\left(2cD_{o}\right)^{1/2}$ by identification with Eq. 5. In our experiments, $D_{o}$ is constant, and $b^{1/4}$ varies slowly, so $A_{o}$ is primarily controlled by the parameter $c$. As a consistency check, $A_{o}$ is plotted as a function of $c^{1/2}$ in Figure 12 in the Appendix D. The linear fit $A_{o}=0.17\times c^{1/2}$ is in good agreement with the independent estimate $b^{1/4}\left(2D_{o}\right)^{1/2}\approx 0.2$. The relation between the variance and the mean of the plastic strain eventually reads $\delta\epsilon_{xy,p}\sim C_{o}\left|\langle\overline{\epsilon_{xy,p}}\rangle\right|$, where $C_{o}=b^{1/4}(2D_{o}c)^{1/2}\left(1-b\right)^{-1}$ remains close to $5$. This enforces $c\propto\frac{(1-b)^{2}}{b^{1/2}}$ to be a monotonically decreasing function of $b$.
We now outline a speculative scenario to explore the functional relationship between $c$ and $b$. As an analogy with a granular medium, we consider a system with a finite number of mesoscopic “sites” that accumulate strain under forward shear $\epsilon_{m}$. Here, $D_{0}$ represents the strain scale of independent deformation sites, matching the variability of forward and backward strains with their mean and the order of magnitude of the largest strain cells shown in Figure 5. During reverse shear, the sites exhibit an average elastic recovery $b\epsilon_{m}$ to which a stochastic term proportional to $c\epsilon_{m}$ is added. $b(\theta_{m})$ quantifies the effective proportion of reversible sites and hence represent a memory parameter while $c(\theta_{m})$ controls how increasing shear amplitude produces random irreversible deformation. Then, the creation of reversible sites through shearing at higher load exhausts the reservoir of sites that can still slip irreversibly. More work is needed to investigate the potential link of our results with the emergence of limit cycles [31].
@@
Plasticity in our system, viewed through the lens of local fluctuations, could be compared to other divided systems. In emulsions [10], fluctuations increase with decreasing packing fraction at controlled strain. In colloidal glasses [11], low packing fraction leads to fewer plastic events, as the lubricated interactions between grains allow for reversibility. Our study suggests distinct signatures of emerging plasticity, shedding light on the effects of applied load and memory.